7 years ago x is 3 times older than y. Now
X is 5 times older tham y. What is the present age of X?

Answer by Alan3354(69443)

7 years ago x is 3 times older than y. Now
X is 5 times older tham y. What is the present age of X?
--------------
1 time older = 2 times as old.
2 times older = 3 times as old.
3 times older = 4 times as old.
4 times older = 5 times as old.
5 times older = 6 times as old.
-----
7 years ago x was 4 times as old as y.
Now X is 6 times as old as y.
What is the present age of X?
---
x-7 = 4*(y-7) = 4y-28
x = 6y
-----
x-7 = 4y-28
x = 4y-21
6y = 4y - 21
2y = -21
y = -10.5 years
x = -63 years